OpenStudy (anonymous):

sqrt2 + sqrt5 _____________ sqrt2 - sqrt5

OpenStudy (anonymous):

\[\sqrt{2} + \sqrt{5} \div \sqrt{2} - \sqrt{5} \]

OpenStudy (anonymous):

oops didn't see your denominator there sqrt2 + sqrt5 Multiply that thing by ----------- sqrt2 + sqrt5

OpenStudy (anonymous):

You get 2 + 2sqrt10 + 5 --------------- 2 - 5 which equals 7 + 2sqrt10 ----------- -3 or -7/3 - 2/3 sqrt10

OpenStudy (anonymous):

oh okay thank you!!

OpenStudy (anonymous):

sure no problem